OpenStudy (sstradic):

MEDAL : A CD tower is shaped like a cylinder and holds CDs with a diameter of 4.75 inches. If the CD tower is 12 inches tall, what is its volume in cubic inches? Use 3.14 for pi. Round your answer to the nearest hundredth. 212.54 cubic inches 231.56 cubic inches 850.16 cubic inches 867.34 cubic inches

Vocaloid (vocaloid):

hints: radius = (1/2) * diameter volume = pi * (radius)^2 * height